OpenStudy (anonymous):

Write the equation in Slope-intercept form and Standard Form. Show all steps. -3+x=y Will award medal.

OpenStudy (anonymous):

slope intercept form is y=mx+b

OpenStudy (anonymous):

do you know what each variable means?

OpenStudy (anonymous):

I think so. I know that slope form will be y=x+-3

OpenStudy (anonymous):

yes, or y=x-3

OpenStudy (anonymous):

do you know what standard form is?

OpenStudy (anonymous):

isn't it y2-y1/x2-x1

OpenStudy (anonymous):

I don't think so. Watch this video. It should help you understand it without me giving you the answer: https://www.youtube.com/watch?v=IJZ9nXYgwqc

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

get it yet?

OpenStudy (anonymous):

Would it be 1x+-1y=-3

OpenStudy (anonymous):

Oh wait I messed up wright there didn't I?

OpenStudy (anonymous):

right*

OpenStudy (anonymous):

don't use 1x or 1y, use x or y instead because 1x is the same as x

OpenStudy (anonymous):

one second...

OpenStudy (anonymous):

I think it would be y-x=-3

OpenStudy (anonymous):

I have to check my answer though

OpenStudy (anonymous):

ok I can wait.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

that's the right answer? @heretohelpalways

OpenStudy (anonymous):

my answer is right: y-x=-3

OpenStudy (anonymous):

I think it is, and tag someone just in case

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

@Nnesha could you check this answer please?

OpenStudy (anonymous):

Ax+By=C is standard form equation

Nnesha (nnesha):

y-x= -3 is same as -x+y = -3 would be great if you multiply/divide both sides by negative one to get the positive Ax

OpenStudy (anonymous):

wouldn't it be x+y=-3?

Nnesha (nnesha):

looks good to me

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Ok thank you both of you guys. Much appreciated

OpenStudy (anonymous):

wait,I can tell you why it isn't x+y=-3

OpenStudy (anonymous):

so when converting this : y=x-3 into standard form, I think of it as getting -3 by itself

Nnesha (nnesha):

i can understand i know what you are going to say

Nnesha (nnesha):

but the standard form is \[\huge\rm Ax+By=C\] where A is positive integer we need x-term to be positive :=)

Nnesha (nnesha):

good job to both of you !! \o^_^o/ great work